The Corata Appliance Manufacturing Corporation manufactures two vacuum cleaners, the Standard and the Super. The following information was gathered about the two products:

                                                                                           Standard               Super

    Budgeted sales in units                                                 2700                       700

    Budgeted selling price                                                   $600                  $1,700

    Budgeted contribution margin per unit                   $500                   $1070

    Actual sales in units                                                      2800                     1400

    Actual selling price                                                        $650                  $1,680

What is the total sales-quantity variance in terms of the contribution margin? (Round intermediary calculations to two decimal places.)

A.

$495,760 favorable

B.

$179,760 favorable

C.

$316,000 favorable

D.

$136,240 favorable
Answer #1

Budgeted sales mix

Standard = 2700/(2700+700) = 0.79

Super = 700/(2700+700) = 0.21

Total sales quantity Variance

= (Actual sales-Budgeted) * budgeted sales mix * budgeted contribution margin

Standard = [(2800+1400)-(2700+700)]*0.79*500

= 316,000

Super = [(2800+1400)-(2700+700)]*0.21*1070

= 179,760

Total = 316,000+179,760

= 495,760 Favourable